-x+2y=6 i don't understand how to graph this
Answer by lwsshak3(11628)

-x+2y=6 i don't understand how to graph this
**
You can always plot points to graph this function, but there is better way to graph the function and help you understand it at the same time.
First, you should recognize this is a first degree equation so the graph will be a straight line.
The standard form of equation for a straight line is y=mx+b, m=slope, b=y-intercept.
Rewrite equation into standard form:
2y=x+6
y=x/2+3
This is an equation of a straight line with slope=1/2 and y-intercept=3
Graphing:
Starting from a point at the y-intercept, go up one unit and 2 units to the right (rise/run) to get the second point. A straight line graph will go thru these two points. To check your solution, plug in an x-value and see if the point falls on the line.